Please help I got them wrong and would like to know what was the right answers
The boxplot below shows salaries for Construction workers and Teachers. Construction Teacher 20 25 30 35 40 45 50 Salary (thousands of S) If a person is making the first quartile salary for a construction worker, they are making more than what percentage of Teachers? They are making more than 75-..., % of Teachers. Question 2. Last Attempt:0 out of 1 Score in Gradebook:0 out of 1
Ages Number of students 15-18 19-22 23-26 27-30 31-34 35-38 2 10 Based on the frequency distribution above, find the cumulative frequency for the class with lower class limit 31 Cumulative Frequency: 12.5 Give your answer rounded to one decimal place Question 8. Last Attempt:0 out of 2 Score in Gradebook:0 out of 2
Answer #1

Sol:

Ages Number of student Cumulative frequency
15-18 2 2
19-22 6 2+6=8
23-26 8 2+6+8=16
27-30 10 2+6+8+10=26
31-34 5 2+6+8+10+5=31
35-38 9 2+6+8+10+5+9=40

based on the table above

31-34 2+6+8+10+5=31

ANSWER:31